# On the feasibility of performing quantum chemistry calculations on quantum computers

### v3

@@ II A criterion for quantum phase estimation (QPE): estimating the orthogonality catastrophe
Figure 4: (a) Sketch of the energy $E(\tau)$ versus imaginary time. The area $\kappa$ under the energy curve directly provides the overlap $\Omega=e^{-\kappa}$. We approximate $\kappa$ with the more easily accessible orange-shaded area $I_{\Omega}$. (b), (c) and (d): Energy $E_{V}$ versus optimization step in VQE simulation, the ”error bar” corresponds to the standard deviation $\sigma_{V}$. (e), (f) and (g): Overlap $\Omega$ and $e^{-I_{\Omega}}$ versus optimization step.
Figure 5: Scaling of the Hartree-Fock energy (a) and the energy variance (b) versus number of atoms in a hydrogen chain. (c): Overlap index $I_{\Omega}$ versus energy error $|E-E_{0}|$ for the variational ansatz of the [51] data set. The dashed line is a linear fit $I_{\Omega}\approx 27.8|E-E_{0}|$.
@@ II.1 Role of the overlap in QPE
QPE starts from a guess input state $|\Psi_{V}\rangle$ and extracts one of the eigenenergies $E_{i}$ of $H$ with a probability $p_{i}=|\langle\Psi_{V}|\Psi_{i}\rangle|^{2}$ while projecting the input state to the corresponding eigenstate $|\Psi_{i}\rangle$ of $H$. In general, one designs $|\Psi_{V}\rangle$ to have the largest overlap with the ground eigenstate $|\Psi_{0}\rangle$ of $H$, so that we will henceforth set $i=0$, and call
$\Omega=|\langle\Psi_{V}|\Psi_{0}\rangle|^{2}$ (18)
the overlap of $|\Psi_{V}\rangle$ with the ground state. $\Omega$ is thus also the success probability of QPE. In practice, it means that QPE needs to be repeated $O(1/\Omega)$ times to find the ground state energy with high probability.
@@
Thus, a key necessary condition for the success of QPE is a large overlap $\Omega\sim 1$ of the input state to the ground state. For QPE to be of practical use, it is therefore necessary to provide the QPE user with a practical means to quantify the overlap in realistic situations where the solution is not known a priori. This is all the more crucial as the state preparation may become exponentially difficult for large systems, i.e. in the very region where QPE is expected to show an advantage over classical or NISQ methods. Indeed, in condensed matter, the overlap between two states, however close they are at the single-body level, is believed to decrease exponentially with system size; this is the orthogonality catastrophe [52]. This phenomenon holds even for states that share very similar energies: two states can be infinitesimally close in energy but physically very distinct. This means in particular that there is no guarantee that a variational state prepared with an algorithm such as VQE will be a good candidate for QPE. For instance, the difference of energy per particle between superconducting aluminum and normal-metal aluminum is extremely tiny $\sim(\Delta/E_{F})^{2}\approx 10^{-8}$ (with $\Delta$ the superconducting gap and $E_{F}$ the Fermi energy), yet the two states behave drastically differently. In quantum chemistry, the situation has been studied in less detail: Tubman et al. [55] looked at small molecules with small static correlations (good targets for classical calculations) and concluded that $\Omega$ could be kept to relatively high values, while [53], which looked at more correlated molecules, gave indication of a relatively fast decay of the overlap. We note that one should be careful with extrapolating calculations made on small molecules: they may be misleadingly optimistic because the variational ansatz may have enough parameters to capture the entire Hilbert space (or a large fraction), in sharp contrast to real situations of interest.
@@ II.2 A criterion for estimating the overlap and the orthogonality catastrophy
Here, we show that the success probability $\Omega$ can actually be estimated in realistic situations where one does not have access to the exact ground state. Let us assume that the initial state $|\Psi_{V}\rangle$ fed to QPE has been obtained, for instance using a variational computation like VQE. Then, we use a theorem proved by one of us in [56] to estimate the overlap $\Omega$ (18) of this state $|\Psi_{V}\rangle$ with the ground state $|\Psi_{0}\rangle$. Let us consider the wavefunction $|\Psi(\tau)\rangle=\frac{1}{\sqrt{Z}}e^{-H\tau}|\Psi_{V}\rangle$, where the factor $Z=\langle\Psi_{V}|e^{-2H\tau}|\Psi_{V}\rangle$ ensures normalization. This wavefunction appears in various techniques (e.g. Diffusion Monte-Carlo or Green function Monte-Carlo) that project the variational wavefunction onto the groundstate using so-called imaginary-time evolution, since $\lim_{\tau\rightarrow\infty}|\Psi(\tau)\rangle=|\Psi_{0}\rangle$. A typical output of these methods is the energy $E(\tau)=\langle\Psi(\tau)|H|\Psi(\tau)\rangle$ as a function of $\tau$ as sketched in Fig. 4 (left panel). The overlap $\Omega$, giving the success probability of QPE, is simply related to the area $\kappa$ under this curve as [56],
[…]
While $\kappa$ is not necessarily easy to compute, we find that a good proxy can be obtained by considering the area $I_{\Omega}$ of the dashed triangle in the left panel of Fig. 4. This area can be calculated from the knowledge of the variational energy $E_{V}=E(\tau=0)$, the energy variance of the variational ansatz $\sigma_{V}^{2}=-\partial_{\tau}E(\tau=0)$, and an estimate (that need not be very accurate) of the ground-state energy $E_{0}$:
$I_{\Omega}\equiv\frac{(E_{V}-E_{0})^{2}}{2\sigma^{2}}.$ (20)
We call $I_{\Omega}$ the ”overlap index” of the variational ansatz $|\Psi_{V}\rangle$. $I_{\Omega}$ provides an estimate of $\Omega$ through the relation
@@
While Eq. (19) is an exact formula, Eq. (21) is an approximation that relies on the energy $E(\tau)$ to smoothly relax to the ground state. When $E(\tau)$ is convex (this has always been the case in previous Green’s function quantum Monte-Carlo calculations performed by the authors), then Eq. (20) gives an exact upper bound $\Omega\leq e^{-I_{\Omega}}$, which is sufficient for our purpose. The associated success criterion for QPE is naturally $\Omega\sim 1$, i.e. $I_{\Omega}\ll 1$. This criterion depends on the energy and variance of the ansatz. It is totally independent on the imaginary-time evolution used for its derivation, and it is much easier to calculate than the actual overlap $\Omega$. Actually, in cases where one would be able to calculate $\Omega$ directly, there would simply be no need for a quantum computer, as one would already have the entire information on the given state.
To corroborate the validity of this overlap estimation, we have performed VQE simulations for several small molecules in simple orbital bases, computing both the variational energy and the variance. For these small systems (up to 8 qubits), the exact ground state energy in the given orbital basis can be calculated as well using exact diagonalization (aka full configuration interaction). This allows to compute the exact overlap $\Omega$ and check that the estimate Eq. (21) holds. On bigger systems, one would rely on estimates for $E_{0}$ such as those currently used in quantum chemistry where one extrapolates from a sequence of increasingly accurate calculations (e.g. CCSD, CCSDT and CCSDTQ calculations). We use the myQLM-fermion package [57], a one-layer UCC ansatz and a minimum basis set (STO-$3$G for H2 and H4 molecules; the $6$-$31$G basis and active space selection to reduce the Hilbert space dimension from $22$ to $4$ for LiH).
The convergence of the results versus optimization steps are shown on the right panels of Figure 4. Note that the ”error bars” in the upper panels stand for the variance of the variational ansatz. In the lower panels, we observe a very good match between the right and left-hand sides of Eq. (21), which shows that the overlap index is a good proxy for $\Omega$. Our results indicate that the joined calculation of the energy $E_{V}$ and the variance $\sigma_{V}^{2}=\langle H^{2}\rangle-E_{V}^{2}$ makes variational calculations much more valuable. Note that in the H4 simulation, which uses $8$ qubits in contrast to H2, and LiH, where the number of qubits is only $4$, the convergence is much slower.
@@
This is the orthogonality catastrophe in the context of variational calculations. To illustrate the above statements, the top panels of Figure 5 show the Hartree-Fock energy (left) and variance (right) of hydrogen chains of up to $N=128$ in the STO-$3$G basis set. Both indeed scale linearly with $N$, as advertised.
@@ III Conclusion
To conclude, we have proposed two criteria, one for VQE (noisy hardware) and one for QPE (fault-tolerant hardware), that are easily accessible and provide necessary conditions for the possibility of doing genuinely relevant chemistry calculations on quantum hardware. Our preliminary estimates imply that this possibility is unlikely with the approaches and technologies that are currently pursued. Decoherence on NISQ devices appears as a very important roadblock for VQE, not to mention shot noise and barren plateau issues. Ways to circumvent the latter two may involve a deeper hybridization between classical and quantum codes, as in [58], where the quantum computer is used only to produce the most relevant electronic configurations (but with no guarantee that a gain with respect to a purely classical selected configuration interaction method would not yield better results), or, in general, in hybrid Monte-Carlo codes (see e.g [59] for a review). As for decoherence, ideas to use noise to one’s advantage have emerged recently [60, 61, 62].
@@ Acknowledgements
We acknowledge funding from the French ANR QPEG and the Plan France 2030 ANR-22-PETQ-0007 “EPIQ”.

### v4

@@ II A criterion for quantum phase estimation (QPE): estimating the orthogonality catastrophe
Figure 4: (a) Sketch of the energy $E(\tau)$ vs imaginary time. The area $\kappa$ under the energy curve directly provides the overlap $\Omega=e^{-\kappa}$. We approximate $\kappa$ with the more easily accessible orange-shaded area $I_{\Omega}$. [(b) - (d)]: Energy $E_{V}$ vs optimization step in VQE simulation, the ”error bar” corresponds to the standard deviation $\sigma_{V}$. [](e) - (g)]: Overlap $\Omega$ and $e^{-I_{\Omega}}$ vs optimization step.
Figure 5: Scaling of the Hartree-Fock energy (a) and the energy variance (b) vs number of atoms in a hydrogen chain. (c): Overlap index $I_{\Omega}$ versus energy error $|E-E_{0}|$ for the variational ansatz of the [67] data set. The dashed line is a linear fit $I_{\Omega}\approx 27.8|E-E_{0}|$.
@@ II.1 Role of the overlap in QPE
QPE starts from a guess input state $|\Phi\rangle$ and extracts one of the eigenenergies $E_{i}$ of $H$ with a probability $p_{i}=|\langle\Phi|\Psi_{i}\rangle|^{2}$ while projecting the input state to the corresponding eigenstate $|\Psi_{i}\rangle$ of $H$. In general, one designs $|\Phi\rangle$ to have the largest overlap with the ground eigenstate $|\Psi_{0}\rangle$ of $H$, so that we will henceforth set $i=0$, and call
$\Omega=|\langle\Phi|\Psi_{0}\rangle|^{2}$ (25)
the overlap of $|\Phi\rangle$ with the ground state. $\Omega$ is thus also the success probability of QPE. In practice, it means that QPE needs to be repeated $O(1/\Omega)$ times to find the ground state energy with high probability (or possibly $O(1/\Omega^{1/2})$ using optimal algorithms [70]).
@@
Thus, a key necessary condition for the success of QPE is a large overlap $\Omega\sim 1$ of the input state to the ground state. For QPE to be of practical use, it is therefore necessary to provide the QPE user with a practical means to quantify the overlap in realistic situations where the solution is not known a priori. This is all the more crucial as the state preparation may become exponentially difficult for large systems, i.e. in the very region where QPE is expected to show an advantage over classical or NISQ methods. Indeed, in condensed matter, the overlap between two states, however close they are at the single-body level, is believed to decrease exponentially with system size; this is the orthogonality catastrophe [68]. This phenomenon holds even for states that share very similar energies: two states can be infinitesimally close in energy but physically very distinct. This means in particular that there is no guarantee that a variational state prepared with an algorithm such as VQE will be a good candidate for QPE. For instance, the difference of energy per particle between superconducting aluminum and normal-metal aluminum is extremely tiny $\sim(\Delta/E_{F})^{2}\approx 10^{-8}$ (with $\Delta$ the superconducting gap and $E_{F}$ the Fermi energy), yet the two states behave drastically differently. In quantum chemistry, the situation has been studied in less detail: [72] looked at small molecules with small static correlations (good targets for classical calculations) and concluded that $\Omega$ could be kept to relatively high values. Similar conclusions were found in [73] for a larger compound. In contrast, [69], which looked at more correlated molecules, gave indication of a relatively fast decay of the overlap. We note that one should be careful with extrapolating calculations made on small molecules: they may be misleadingly optimistic because the variational ansatz may have enough parameters to capture the entire Hilbert space (or a large fraction), in sharp contrast to real situations of interest. The overlap problem has lead some authors to recently consider some mitigation strategies that rely on heavy classical calculations to provide an ansatz with high enough fidelity. One possibility would be to use the results of a DMRG calculation since the corresponding matrix product states can be associated to a quantum circuits and then directly encoded in a quantum computer [74]. In a different context, it had already been shown that, under some circumstances, a quantum circuit improves on a classical result (Corollary E.1. in [75]). We note, however, that in such a scheme, one remains constrained by the limitations of the classical approach so the quantum computer could at best give some extra precision for molecules that are already accessible to classical calculations. Whether this could be sufficient to bring an insufficient guess to chemical accuracy remains to be shown.
@@ II.2 A criterion for estimating the overlap and the orthogonality catastrophy
Here, we show that the success probability $\Omega$ can actually be estimated in realistic situations where one does not have access to the exact ground state. Let us assume that the initial state $|\Phi\rangle$ fed to QPE has been obtained, for instance using a variational computation like VQE. However note that our demonstration is general, valid for any given state preparation method. Then, we use a theorem proved by one of us in [76] to estimate the overlap $\Omega$ (25) of this state $|\Phi\rangle$ with the ground state $|\Psi_{0}\rangle$. Let us consider the wavefunction $|\Psi(\tau)\rangle=\frac{1}{\sqrt{Z}}e^{-H\tau}|\Phi\rangle$, where the factor $Z=\langle\Phi|e^{-2H\tau}|\Phi\rangle$ ensures normalization. This wavefunction appears in various techniques (e.g. Diffusion Monte Carlo or Green function Monte Carlo) that project the initial wavefunction onto the ground state using so-called imaginary-time evolution, since $\lim_{\tau\rightarrow\infty}|\Psi(\tau)\rangle=|\Psi_{0}\rangle$. A typical output of these methods is the energy $E(\tau)=\langle\Psi(\tau)|H|\Psi(\tau)\rangle$ as a function of $\tau$ as sketched in Fig. 4 (left panel). The overlap $\Omega$, giving the success probability of QPE, is simply related to the area $\kappa$ under this curve as [76],
$\Omega=\exp(-\kappa),\mathrm{~with~}\kappa=\int\limits_{0}^{\infty}d\tau(E(\tau)-E_{0}).$ (26)
While $\kappa$ is not necessarily easy to compute, we find that a good proxy can be obtained by considering the area $I_{\Omega}$ of the dashed triangle in the left panel of Fig. 4. This area can be calculated from the knowledge of the initial energy $E_{\Phi}=E(\tau=0)$, the energy variance of the preparation ansatz $\sigma_{\Phi}^{2}=-\partial_{\tau}E(\tau=0)$, and an estimate (that need not be very accurate) of the ground-state energy $E_{0}$:
$I_{\Omega}\equiv\frac{(E_{\Phi}-E_{0})^{2}}{2\sigma_{\Phi}^{2}}.$ (27)
We call $I_{\Omega}$ the ”overlap index” of the ansatz $|\Phi\rangle$. $I_{\Omega}$ provides an estimate of $\Omega$ through the relation
@@
While Eq. (26) is an exact formula, Eq. (28) is an approximation that relies on the energy $E(\tau)$ to smoothly relax to the ground state. When $E(\tau)$ is convex (this has always been the case in previous Green’s function quantum Monte Carlo calculations performed by the authors), then Eq. (27) gives an exact upper bound $\Omega\leq e^{-I_{\Omega}}$, which is sufficient for our purpose. The associated success criterion for QPE is naturally $\Omega\sim 1$, i.e. $I_{\Omega}\ll 1$. This criterion depends on the energy and variance of the ansatz. It is totally independent on the imaginary-time evolution used for its derivation, and it is much easier to calculate than the actual overlap $\Omega$. Actually, in cases where one would be able to calculate $\Omega$ directly, there would simply be no need for a quantum computer, as one would already have the entire information on the given state.
To corroborate the validity of this overlap estimation, we have performed VQE simulations for several small molecules in simple orbital bases, computing both the variational energy and the variance. For these small systems (up to 8 qubits), the exact ground state energy in the given orbital basis can be calculated as well using exact diagonalization (aka full configuration interaction). This allows to compute the exact overlap $\Omega$ and check that the estimate Eq. (28) holds. On bigger systems, one would rely on estimates for $E_{0}$ such as those currently used in quantum chemistry where one extrapolates from a sequence of increasingly accurate calculations (e.g. CCSD, CCSDT and CCSDTQ calculations). We use the myQLM-fermion package [77], a one-layer UCC ansatz and a minimum basis set (STO-33G for H2 and H4 molecules; the 66-3131G basis and active space selection to reduce the Hilbert space dimension from $22$ to $4$ for LiH).
The convergence of the results versus optimization steps is shown on the right panels of Figure 4. Note that the ”error bars” in the upper panels stand for the variance of the variational ansatz. In the lower panels, we observe a very good match between the right and left-hand sides of Eq. (28), which shows that the overlap index is a good proxy for $\Omega$. Our results indicate that the joined calculation of the energy $E_{V}$ and the variance $\sigma_{V}^{2}=\langle H^{2}\rangle-E_{V}^{2}$ makes variational calculations much more valuable. Note that in the H4 simulation, which uses $8$ qubits in contrast to H2, and LiH, where the number of qubits is only $4$, the convergence is much slower.
@@
This is the orthogonality catastrophe in the context of variational calculations. To illustrate the above statements, the top panels of Figure 5 show the Hartree-Fock energy (left) and variance (right) of hydrogen chains of up to $N=128$ in the STO-33G basis set. Both indeed scale linearly with $N$, as advertised.
@@
We note that this is a generic problem of phase estimation algorithms: even more recent version of the algorithm assume an input state with a large overlap (see e.g [78], which even requires a lower bound on this overlap, and [79], which does not).
@@ III Conclusion
To conclude, we have proposed two criteria, one for VQE (noisy hardware) and one for QPE (fault-tolerant hardware), that are easily accessible and provide necessary conditions for the possibility of doing genuinely relevant chemistry calculations on quantum hardware. Our preliminary estimates imply that this possibility is unlikely with the approaches and technologies that are currently pursued. Decoherence on NISQ devices appears as a very important roadblock for VQE, not to mention shot noise and barren plateau issues. There are multiple proposals to circumvent the latter two problems. One may involve a deeper hybridization between classical and quantum codes, as in [57, 58], where the quantum computer is used only to produce the most relevant electronic configurations, or, in general, in hybrid Monte Carlo codes (see e.g [80] for a review). There is, however, no guarantee that a purely classical selected configuration interaction method would not yield better results. As for decoherence, ideas to use noise to one’s advantage have emerged recently [81, 82, 83].
@@ Acknowledgements
We acknowledge funding from the French ANR QPEG and the ANR EPIQ. XW acknowledges funding from the PEPR EQUBITFLY, the ANR DADI, the ANR TKONDO and the CEA-FZJ French-German project AIDAS.